I.In each of Problems 23 through 30, use the method of reduction of order to find a second solution of the given differential equation.

1.  t2y″ − 4ty′ + 6y = 0, y1(t) = t2

2.  xy″ − y′ + 4x3y = 0, x > 0; y1(x) = sin x2


1
Expert's answer
2022-01-04T12:16:59-0500

1.

We seek a solution of the form "y(t) = v(t)y_1(t)." This gives us


"y(t)=v(t)t^2"

"y'(t)=v't^2+2tv"

"y''(t)=v''t^2+4tv'+2v"

Substitute


"t^4v''+4t^3v'+2t^2v\u2212 4t^3v'-8t^2v+6t^2v = 0"

"t^4v'' = 0"

"v''=0"

Integrate


"v'=c_1"

Integrate

"v=c_1t+c_2"

Therefore, the general solution is


"y(t)=(c_1t+c_2)t^2"

"y(t)=c_1t^3+c_2t^2"

The second solution is "y_2(t)=t^3."


2.

We seek a solution of the form "y(x) = v(x)y_1(x)." This gives us


"y(x)=v(x)\sin(x^2)"

"y'(x)=v'\sin(x^2)+2x\cos(x^2)v"

"y''(x)=v''\sin(x^2)+2x\cos(x^2)v'+2\cos(x^2)v"

"-4x^2\sin(x^2)v+2x\cos(x^2)v'"

Substitute


"x\sin(x^2)v''+4x^2\cos(x^2)v'+2x\cos(x^2)v"

"-4x^3\sin(x^2)v-\sin(x^2)v'-2x\cos(x^2)v"

"+4x^3\sin(x^2)v(x)=0"

"x\sin(x^2)v''+4x^2\cos(x^2)v'-\sin(x^2)v'=0"

"x\sin(x^2)v''=(-4x^2\cos(x^2)+\sin(x^2))v'"

"\dfrac{v''}{v'}=-4x\cos(x^2)+\dfrac{\sin(x^2)}{x}""d(\ln(v'))=(-4x\dfrac{\cos(x^2)}{\sin(x^2)}+\dfrac{1}{x})dx"

Integrate


"\int d(\ln(v'))=\int (-4x\dfrac{\cos(x^2)}{\sin(x^2)}+\dfrac{1}{x})dx"


Find


"-\int4x\dfrac{\cos(x^2)}{\sin(x^2)}dx"

"u=\sin(x^2), du=2x\cos(x^2)dx"

"-\int4x\dfrac{\cos(x^2)}{\sin(x^2)}dx=-2\int \dfrac{du}{u}=-2\ln(u)+\ln C_1"

"=-2\ln(\sin(x^2))+\ln C_1"

"\ln(v')=\ln(\sin^{-2}(x^2))+\ln(x)+\ln C_1"

"v'=\dfrac{C_1 x}{\sin^2(x^2)}"

Integrate


"v=\int \dfrac{C_1 x}{\sin^2(x^2)}dx"

Find


"\int\dfrac{C_1 x}{\sin^2(x^2)}dx"

"u=x^2, du=2xdx"

"\int\dfrac{C_1 x}{\sin^2(x^2)}dx=\dfrac{C_1}{2}\int\dfrac{1}{\sin^2(u)}du"

"=C_2\cot(u)+C_3=C_2\cot(x^2)+C_3"

"v(x)=C_2\cot(x^2)+C_3"

Then


"y(x)=(C_2\cot(x^2)+C_3)\sin(x^2)"

"=C_2\cos(x^2)+C_3\sin(x^2)"

Therefore, the general solution is


"y(x)=C_2\cos(x^2)+C_3\sin(x^2)"

The second solution is "y_2(x)=\cos(x^2)."